OpenStudy (anonymous):

0.3(x+0.2) + 0.4x=5.5

OpenStudy (anonymous):

Simplifying 0.3(x + 0.2) + 0.4x = 5.5 Reorder the terms: 0.3(0.2 + x) + 0.4x = 5.5 (0.2 * 0.3 + x * 0.3) + 0.4x = 5.5 (0.06 + 0.3x) + 0.4x = 5.5 Combine like terms: 0.3x + 0.4x = 0.7x 0.06 + 0.7x = 5.5 Solving 0.06 + 0.7x = 5.5 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-0.06' to each side of the equation. 0.06 + -0.06 + 0.7x = 5.5 + -0.06 Combine like terms: 0.06 + -0.06 = 0.00 0.00 + 0.7x = 5.5 + -0.06 0.7x = 5.5 + -0.06 Combine like terms: 5.5 + -0.06 = 5.44 0.7x = 5.44 Divide each side by '0.7'. x = 7.771428571 Simplifying x = 7.771428571

OpenStudy (anonymous):

wrong answer.